# The weak Extension Principle

### v1

@@ Abstract.
We prove a rigidity result for maps between Čech-Stone remainders under fairly mild forcing axioms.
@@ 1. Introduction
A natural way to construct a continuous map between Čech–Stone remainders is to consider the restriction of the compactification of a continuous map between the underlying spaces (perhaps forgetting about compact sets). To be precise, if $X$ and $Y$ are locally compact noncompact spaces, and $V_{X}\subseteq X$ is an open set with compact closure [^1] The need of $V_{X}$ is justified, for example, by the fact that the spaces $\mathbb{R}^{*}$ and $((-\infty,0]\cup[1,\infty))^{*}$ are homeomorphic, but no continuous function $\mathbb{R}\to(-\infty,0]\cup[1,\infty)$ can induce a homeomorphism between them., a proper continuous map $X\setminus V_{X}\to Y$ extends to a map $\beta(X\setminus V_{X})\to\beta Y$ whose restriction to $X^{*}$ gives a continuous map $X^{*}\to Y^{*}$. Does this construction recover all continuous maps between Čech–Stone remainders, at least without appealing to unnatural back-and-forth constructions? This question was first analyzed for the simplest space among all, $\mathbb{N}$ with the discrete topology, and it was given a negative answer: in [3], Dow constructed (in $\operatorname{{\mathsf{ZFC}}}$) a nontrivial copy of ${\mathbb{N}}^{*}$ inside ${\mathbb{N}}^{*}$ which is nowhere dense, and therefore cannot come from a map ${\mathbb{N}}\to{\mathbb{N}}$ as above (further considerations on nontrivial copies of ${\mathbb{N}}^{*}$ were recently made in [1] and [2]). Conjecturally, this is all there is in $\operatorname{{\mathsf{ZFC}}}$, and under some reasonably weak forcing axioms, all maps between Čech–Stone remainders arise as a blend of two maps as above. This intuition was formalised by Farah, who in [5, §4] introduced the weak Extension Principle. This principle was stated originally for pairs of zero-dimensional spaces; the following adapts it to the class of locally compact noncompact second countable spaces:
@@
By $\operatorname{{\mathsf{wEP}}}$ we denote the statement “$\operatorname{{\mathsf{wEP}}}(X,Y)$ holds whenever $X$ and $Y$ are locally compact, noncompact second countable spaces”.
The weak Extension Principle cannot be a consequence of $\operatorname{{\mathsf{ZFC}}}$, as witnessed by the possible existence of nontrivial autohomeomorphisms of $\mathbb{N}^{*}$ under the Continuum Hypothesis $\operatorname{\mathsf{CH}}$ ([16]). More in depth, by a classical result of Parovičenko ([15], $\operatorname{\mathsf{CH}}$ implies that if $X$ is a zero-dimensional locally compact noncompact second countable space, then $X^{*}$ is homeomorphic to ${\mathbb{N}}^{*}$. All of these nontrivial homeomorphisms witness the failure of $\operatorname{{\mathsf{wEP}}}$. Further, again under $\operatorname{\mathsf{CH}}$, all Čech–Stone remainders (of second countable spaces) are continuous images of ${\mathbb{N}}^{*}$, and all connected ones are continuous images of $[0,1)^{*}$. A simple counting argument that shows most of these maps cannot arise from continuous maps from ${\mathbb{N}}$ (or $[0,1)$) to the locally compact space of interest. In general, $\operatorname{\mathsf{CH}}$ and back-and-forth constructions make essentially impossible to classify maps between Čech–Stone remainders.
Is it then possible that $\operatorname{{\mathsf{wEP}}}$ is consistent? Instances of $\operatorname{{\mathsf{wEP}}}$ were analysed first for $X=Y={\mathbb{N}}$, and more generally to study maps between finite powers of $\mathbb{N}^{*}$ in different models of set theory (see [12, 7, 13, 17, 4] and [18], or [8] for more related questions). Farah, while formally introducing the $\operatorname{{\mathsf{wEP}}}$, showed under some fairly mild forcing axioms (which are, Todorcevic̀’s $\operatorname{{\mathsf{OCA}}}$ and Martin’s Axioms $\operatorname{{\mathsf{MA}_{\aleph_{1}}}}$), $\operatorname{{\mathsf{wEP}}}(X,Y)$ holds for pairs of countable locally compact noncompact spaces. Further results were obtained in [11] and [19], focusing on just homeomorphisms instead of general continuous maps. For a lengthy and detailed treatment on how forcing axioms impact the structure of homomorphisms of quotient structures, see [10].
@@
As a consequence, we show that the thesis of [6, Theorem 5.1] holds (under $\operatorname{{\mathsf{OCA}}}$ and $\operatorname{{\mathsf{MA}_{\aleph_{1}}}}$) for all pairs of locally compact noncompact second countable spaces: if $\kappa$ and $\lambda$ are infinite cardinals with $\kappa<\lambda$ then there is no continuous surjection $(X^{*})^{\kappa}\to(Y^{*})^{\lambda}$.
The proof of our main result proceeds as follows: first, we use a result of Farah ([6]) on maps between powers of Čech–Stone remainders (more generally, $\beta{\mathbb{N}}$-spaces) to simplify the situation and show that it is enough to show $\operatorname{{\mathsf{wEP}}}$ for maps where $d=\ell=1$. Secondly, we rely on Gel’fand’s duality, and study -homomorphisms between coronas of abelian $\mathrm{C^{*}}$-algebras. We build on a strong lifting theorem proved in [19] (see Proposition 2.3) which gives that, under forcing axioms, -homomorphisms between abelian corona $\mathrm{C^{*}}$-algebras have well-behaved liftings on certain blocked subalgebras. The main content of our proof is to glue and analyse such liftings. This approach is fairly different from the one of [5, §4], where Boolean algebras of clopen sets were used. Once the dimension increases, Stone duality is replaced by Gel’fand’s one, as clopen sets do not recover completely the topology in this case.
@@ Acknowledgements
This research originated from DY’s Master thesis, written under the guidance of AV at Université Paris Cité. In this period, DY was supported by an INSMI (Institut National des Sciences Mathématiques et de leurs Interactions) grant for master students. The authors would like to thank I. Farah for useful conversations.
@@ 2. Proof of the main result
This main section is fully dedicated to the proof of Theorem 1.2. As mentioned, our first step is to allow us to only consider maps between Čech–Stone remainders. This is due to a powerful result of Farah. Some context: If $d\geq 1$, a function between topological spaces $F\colon X^{d}\to Y$ is said to depend on at most one coordinate on $U\subseteq X^{d}$ if there exists $j\leq d$ and $H\colon X\to Y$ such that $F\restriction U=H\restriction\pi_{j}(U)$, $\pi_{j}\colon X^{d}\to X$ being the projection on the $j$-th coordinate. A rectangle in $X^{d}$ is a set of the form $\prod_{i\leq d}A_{i}$, where $A_{i}\subseteq X$. The following is a reformulation of [6, Theorem 3] [^2] Theorem 3 in [6] works even if $d$ is an infinite cardinal. Further, the codomain space $Y$ needs to be a $\beta{\mathbb{N}}$-space, and not necessarily a remainder space..
@@
First of all, we show that we can assume $\ell=1$. For this, assume the $\operatorname{{\mathsf{wEP}}}$ holds for all maps $(X^{*})^{d}\to Y^{*}$ (where $X$ and $Y$ satisfy the appropriate hypotheses, and $d\geq 1$ is arbitrary). Let $F\colon(X^{*})^{d}\to(Y^{*})^{\ell}$ be a continuous map, and, for $i\leq\ell$, compose $F$ with the projection on the $i$-th coordinate maps. This gives continuous maps $F_{i}=\pi_{i}\circ F\colon(X^{*})^{d}\to Y^{*}$. Applying the weak Extension principle to these maps, we get, for $i\leq\ell$, a clopen decomposition $(X^{*})^{d}=U_{0}^{i}\cup U_{1}^{i}$ and an open set with compact closure $V_{X}^{i}$ such that $F_{i}[U_{0}^{i}]$ is nowhere dense and there is a continuous $G_{i}\colon(\beta X\setminus V_{X}^{i})^{d}\to\beta Y$ which restricts to $F_{i}$ on $U_{1}^{i}$. The sets $U_{0}=\bigcup_{i}U_{0}^{i}$, $U_{1}=\bigcap_{i}U_{1}^{i}$, $V_{X}=\bigcup_{i}V_{X}^{i}$ and the map $G\colon(\beta X\setminus V_{X}^{i})^{d}\to(\beta Y)^{\ell}$ sending $x$ to $(G_{1}(x),\ldots,G_{\ell}(x))$ witness that $F$ satisfies the $\operatorname{{\mathsf{wEP}}}$.
@@
From now on, we only deal with continuous maps between Čech–Stone remainders (i.e., we assume that $d=\ell=1$ in the definition of the $\operatorname{{\mathsf{wEP}}}$). We approach these via Gel’fand’s duality, by looking at the $\mathrm{C^{*}}$-algebras $C(X^{*})$ and $C(Y^{*})$ and at -homomorphisms between them. The structure of such -homomorphisms and their behaviour under forcing axioms was extensively studied in [14] and [19], after the use of forcing axioms to analyse massive quotients of $\mathrm{C^{*}}$-algebras was popularised by the seminal [9].
@@
Since our spaces are second countable and locally compact, they are $\sigma$-compact (alternatively, the $\mathrm{C^{*}}$-algebras $C_{0}(X)$ and $C_{0}(Y)$ are $\sigma$-unital). The following notation is set up to match the hypotheses of [19, Theorem 4.3]. We let
@@
The letters $e$ and $o$ are for even and odd partition. We have the following properties:
@@
Recall, a -homomorphism between nonunital $\mathrm{C^{*}}$-algebras $\psi\colon A\to B$ is nondegenerate if the ideal generated by the image of $\psi$ is dense (alternatively, if an approximate unit for $A$ is sent to an approximate unit of $B$). In the abelian setting, Gel’fand’s duality asserts that nondegenerate -homomorphisms are dual to proper continuous maps. (Recall, a map between topological spaces is proper if inverse images of compact subsets are compact.)
The following is [19, Theorem 4.3] (see also §4.1 therein).
@@
    For $i\in\{e,o\}$ and $n\neq m$, $\overline{V}_{n}^{i}\cap\overline{V}_{m}^{i}=\emptyset$, and
@@
    $\Psi^{i}=\sum_{n}\psi_{n}^{i}\colon\prod_{n\geq\bar{n}}C_{0}(U_{n}^{i})\to\prod_{n\geq\bar{n}}C_{0}(V_{n}^{i})$
@@
First, we show that $V_{1}$ is, modulo compact, clopen: Since
$\overline{U}_{n}^{e}\subseteq U_{n-1}^{o}\cup U_{n}^{e}\cup U_{n}^{o}\text{ and }\overline{U}_{n}^{o}\subseteq U_{n}^{e}\cup U_{n}^{o}\cup U_{n+1}^{e}$
for all $n$, the same applies to all of the $V_{n}^{e}$ and $V_{n}^{o}$ (when $n\geq\bar{n}$). In particular $\overline{V}_{1}\setminus V_{1}$ is contained in the compact set $\overline{V}_{\bar{n}}^{e}$, i.e., $V_{1}$ is, modulo compact, clopen in $X$ [^3] In case the spaces of interest can be written as a union of pairwise disjoint compact open sets, $V_{1}$ can be chosen to be clopen. As the continuous proper maps $\gamma_{n}^{i}$ agree on their common domains, the function
@@
is a well-defined continuous function. Furthermore, $\Gamma$ is proper: if $K\subseteq W$ is compact, then there exists $\bar{n}^{\prime}$ such that $K\subseteq\bigcup_{\bar{n}\leq n\leq\bar{n}^{\prime}}V_{n}^{e}\cup V_{n}^{o}$, hence $\Gamma^{-1}[K]$ is a closed subset of $\bigcup_{\bar{n}\leq n\leq\bar{n}^{\prime}}U_{n}^{e}\cup U_{n}^{o}$. As the latter has compact closure, $\Gamma^{-1}[K]$ is compact. Since $\Gamma$ is proper, it induces a -homomorphism
@@
Since every function $f\in C_{b}(Y)$ can be written (modulo $C_{0}(X)$) as a sum $f=f_{e}+f_{o}$ where, for $i\in\{e,o\}$, $f_{i}\in\prod_{n\geq\bar{n}}C_{0}(U_{n}^{i})$, it is enough to show $\Psi$ lifts $\Phi_{1}$ on $\prod_{n\geq\bar{n}}C_{0}(U_{n}^{i})$, again for $i\in\{e,o\}$. For convenience, let $i=e$, and fix mutually orthogonal positive contractions $g_{n}\in C_{b}(X)$ such that for all $n\geq\bar{n}$
@@
Let $g_{n}=0$ if $n<\bar{n}$. Such functions exist as distinct $V_{n}^{e}$s have disjoint closures. For $S\subseteq{\mathbb{N}}$, let $g_{S}=\sum_{n\in S}g_{n}$.
@@
(2.1) if ($\text{if }(x\in V_{1}\setminus K\text{ and }|\tilde{\Phi}(f_{S})(x)|>\varepsilon)\text{ then }x\in\bigcup_{n\in S}V_{n}^{e}.$
This gives that $g_{S}\tilde{\Phi}(f_{S})=_{C_{0}(X)}\chi_{V_{1}}\tilde{\Phi}(f_{S})$. As $S$ is arbitrary, the same reasoning applied to to ${\mathbb{N}}\setminus S$, gives that $g_{S}\tilde{\Phi}(f_{{\mathbb{N}}\setminus S})\in C_{0}(X)$. Altogether, we get that

### v2

@@ Abstract.
We prove a rigidity result for maps between Čech–Stone remainders under fairly mild forcing axioms.
@@ 1. Introduction
A natural way to construct a continuous map between Čech–Stone remainders is to consider the restriction of the compactification of a continuous map between the underlying spaces (perhaps forgetting about compact sets). To be precise, if $X$ and $Y$ are locally compact noncompact spaces, and $V_{X}\subseteq X$ is an open set with compact closure [^1] The need of $V_{X}$ is justified, for example, by the fact that the spaces $\mathbb{R}^{*}$ and $((-\infty,0]\cup[1,\infty))^{*}$ are homeomorphic, but no continuous function $\mathbb{R}\to(-\infty,0]\cup[1,\infty)$ can induce a homeomorphism between them., a proper continuous map $X\setminus V_{X}\to Y$ extends to a map $\beta(X\setminus V_{X})\to\beta Y$ whose restriction to $X^{*}$ gives a continuous map $X^{*}\to Y^{*}$. Does this construction recover all continuous maps between Čech–Stone remainders, at least without appealing to unnatural back-and-forth constructions? This question was first analyzed for the simplest space among all, $\mathbb{N}$ with the discrete topology, and it was given a negative answer: in [dow2014non], Dow constructed (in $\operatorname{{\mathsf{ZFC}}}$) a nontrivial copy of $\mathbb{N}$$^{*}$ inside $\mathbb{N}$$^{*}$ which is nowhere dense, and therefore cannot come from a map $\mathbb{N}$$\to{\mathbb{N}}$ as above (further considerations on nontrivial copies of $\mathbb{N}$$^{*}$ were recently made in [dow2024autohomeomorphismspreimagesmathbbn] and [dow2024nontrivialcopiesn]). Conjecturally, this is all there is in $\operatorname{{\mathsf{ZFC}}}$, and under some reasonably weak forcing axioms, all maps between Čech–Stone remainders arise as a blend of two maps as above. This intuition was formalised by Farah, who in [Fa:AQ, §4] introduced the weak Extension Principle. This principle was stated originally for pairs of zero-dimensional spaces; the following adapts it to the class of locally compact noncompact second countable spaces:
@@
By $\operatorname{{\mathsf{wEP}}}$ we denote the statement $\operatorname{{\mathsf{wEP}}}(X,Y)$ holds whenever $X$ and $Y$ are locally compact, noncompact second countable spaces”.
The weak Extension Principle cannot be a consequence of $\operatorname{{\mathsf{ZFC}}}$, as witnessed by the existence of nontrivial autohomeomorphisms of $\mathbb{N}^{*}$ under the Continuum Hypothesis $\operatorname{\mathsf{CH}}$ ([Ru]). More in depth, by a classical result of Parovičenko ([Pa:Universal], $\operatorname{\mathsf{CH}}$ implies that if $X$ is a zero-dimensional locally compact noncompact second countable space, then $X^{*}$ is homeomorphic to $\mathbb{N}$$^{*}$. All of these nontrivial homeomorphisms witness the failure of $\operatorname{{\mathsf{wEP}}}$. Further, again under $\operatorname{\mathsf{CH}}$, all Čech–Stone remainders (of second countable spaces) are continuous images of $\mathbb{N}$$^{*}$, and all connected ones are continuous images of $0,1)^{*}$. A simple counting argument shows most of these maps cannot arise from continuous maps from $\mathbb{N}$} (or $[0,1)$) to the locally compact space of interest. In general, $\operatorname{\mathsf{CH}}$ and back-and-forth constructions make essentially impossible to classify maps between Čech–Stone remainders.
Is it then possible that $\operatorname{{\mathsf{wEP}}}$ is consistent? Instances of $\operatorname{{\mathsf{wEP}}}$ were analysed first for $X=Y={\mathbb{N}}$, and more generally to study maps between finite powers of $\mathbb{N}^{*}$ in different models of set theory (see [Just:Omega^n, Fa:Powers, Just:WAT, vD:Prime, DoHa:Images] and [vanMill.Three], or [Fa:Fourth] for more related questions). Farah, while formally introducing the $\operatorname{{\mathsf{wEP}}}$, showed under some fairly mild forcing axioms (which are, Todorcevic̀’s $\operatorname{{\mathsf{OCA}}}$ and Martin’s Axioms $\operatorname{{\mathsf{MA}_{\aleph_{1}}}}$), $\operatorname{{\mathsf{wEP}}}(X,Y)$ holds for pairs of countable locally compact noncompact spaces. Further results were obtained in [FaMcK:Homeomorphisms] and [vignati2018rigidity], focusing on just homeomorphisms instead of general continuous maps. For a lengthy and detailed treatment on how forcing axioms impact the structure of homomorphisms between quotient structures, see [farah2022corona].
@@
As a consequence, we show that the thesis of [Fa:Dimension, Theorem 5.1] holds (under $\operatorname{{\mathsf{OCA}}}$ and $\operatorname{{\mathsf{MA}_{\aleph_{1}}}}$) for all pairs of locally compact noncompact second countable spaces: if $\kappa$ and $\lambda$ are cardinals with $\kappa<\lambda$ then there is no continuous surjection $(X^{*})^{\kappa}\to(Y^{*})^{\lambda}$. This result does not hold under $\operatorname{\mathsf{CH}}$, as under this assumption $\mathbb{N}$$^{*}$ continuously surjects onto every compact topological space of weight $\leq\mathfrak{c}$, and specifically onto every finite power of every Čech–Stone remainder of a locally compact second countable topological space, while $0,1)^{*}$ continuously surjects onto every finite power of every such connected Čech–Stone remainder.
The proof of our main result proceeds as follows: first, we use a result of Farah ([Fa:Dimension]) on maps between powers of Čech–Stone remainders (more generally, $\beta{\mathbb{N}}$-spaces) to simplify the situation and show that it is enough to show $\operatorname{{\mathsf{wEP}}}$ for maps where $d=\ell=1$ (Proposition 2.2). Secondly, we rely on Gel’fand’s duality, and study -homomorphisms between coronas of abelian $\mathrm{C^{*}}$-algebras. We build on a strong lifting theorem proved in [vignati2018rigidity] (see Proposition 2.3) which gives that, under forcing axioms, -homomorphisms between abelian corona $\mathrm{C^{*}}$-algebras have well-behaved liftings on certain blocked subalgebras. The main content of our proof is to glue and analyse such liftings. This approach is fairly different from the one of [Fa:AQ, §4], where Boolean algebras of clopen sets were used. Once the dimension increases, Stone duality is replaced by Gel’fand’s one, as clopen sets do not recover completely the topology in this case.
For all background material on $\mathrm{C^{*}}$-algebras theory we refer the reader to [Black:Operator], or to [Fa:STCstar] for the connections between $\mathrm{C^{*}}$-algebras and set theory. For notions in topology, we refer to the standard book of Munkres ([Munkres.Topology]) and to [GillManJerison.Rings] more specifically for Čech–Stone compactifications and remainders.
@@ Acknowledgements
The authors would like to thank I. Farah for useful conversations, and the anonymous referee for helpful comments.
Funding
This research originated from DY’s Master thesis, written under the guidance of AV at Université Paris Cité. In this period, DY was supported by an INSMI (Institut National des Sciences Mathématiques et de leurs Interactions) grant for master students.
@@ 2. Proof of the main result
This main section is fully dedicated to the proof of Theorem 1.2. As mentioned, our first step is to allow us to only consider maps between Čech–Stone remainders. This is due to a powerful result of Farah. Some context: If $d\geq 1$, a function between topological spaces $F\colon X^{d}\to Y$ is said to depend on at most one coordinate on $U\subseteq X^{d}$ if there exists $j\leq d$ and $H\colon X\to Y$ such that $F\restriction U=H\restriction\pi_{j}(U)$, $\pi_{j}\colon X^{d}\to X$ being the projection on the $j$-th coordinate. A rectangle in $X^{d}$ is a set of the form $\prod_{i\leq d}A_{i}$, where $A_{i}\subseteq X$. The following is a reformulation of [Fa:Dimension, Theorem 3] [^2] Theorem 3 in [Fa:Dimension] works even if $d$ is an infinite cardinal. Further, the codomain space $Y$ needs to be a $\beta{\mathbb{N}}$-space, and not necessarily a remainder space..
@@
The statement of the proposition asserts that if we can prove the weakening of the $\operatorname{{\mathsf{wEP}}}$ where we only consider maps between Čech–Stone remainders, i.e. $d=\ell=1$, then the $\operatorname{{\mathsf{wEP}}}$ holds.
At first, we show that we can assume $\ell=1$. For this, assume the $\operatorname{{\mathsf{wEP}}}$ holds for all maps $(X^{*})^{d}\to Y^{*}$ (where $X$ and $Y$ satisfy the appropriate hypotheses, and $d\geq 1$ is arbitrary). Let $F\colon(X^{*})^{d}\to(Y^{*})^{\ell}$ be a continuous map, and, for $i\leq\ell$, compose $F$ with the projection on the $i$-th coordinate maps. This gives continuous maps $F_{i}=\pi_{i}\circ F\colon(X^{*})^{d}\to Y^{*}$. Applying the weak Extension principle to these maps, we get, for $i\leq\ell$, a clopen decomposition $(X^{*})^{d}=U_{0}^{i}\cup U_{1}^{i}$ and an open set with compact closure $V_{X}^{i}$ such that $F_{i}[U_{0}^{i}]$ is nowhere dense and there is a continuous $G_{i}\colon(\beta X\setminus V_{X}^{i})^{d}\to\beta Y$ which restricts to $F_{i}$ on $U_{1}^{i}$. The sets $U_{0}=\bigcup_{i}U_{0}^{i}$, $U_{1}=\bigcap_{i}U_{1}^{i}$, $V_{X}=\bigcup_{i}V_{X}^{i}$ and the map $G\colon(\beta X\setminus V_{X}^{i})^{d}\to(\beta Y)^{\ell}$ sending $x$ to $(G_{1}(x),\ldots,G_{\ell}(x))$ witness that $F$ satisfies the $\operatorname{{\mathsf{wEP}}}$.
@@
From now on, we only deal with continuous maps between Čech–Stone remainders (i.e., we assume that $d=\ell=1$ in the definition of the $\operatorname{{\mathsf{wEP}}}$). We approach these via Gel’fand’s duality, by looking at the $\mathrm{C^{*}}$-algebras $C(X^{*})$ and $C(Y^{*})$ and at -homomorphisms between them. In this article, we only consider abelian $\mathrm{C^{*}}$-algebras, i.e., algebras of continuous functions on locally compact topological spaces where the topology is induced by the $\sup$-norm and operations are defined pointwise. For an introduction and basic terminology, see [Black:Operator, II.2].
Algebras of the form $C(X^{*})$ are typical examples of corona $\mathrm{C^{*}}$-algebras ([Black:Operator, II.7] or [Fa:STCstar, §13]). The structure of -homomorphisms between corona algebras and their behaviour under forcing axioms was extensively studied in [mckenney2018forcing] and [vignati2018rigidity], after the use of forcing axioms to analyse massive quotients of $\mathrm{C^{*}}$-algebras was popularised by the seminal [Fa:All].
@@
Since our spaces are second countable and locally compact, they are $\sigma$-compact (alternatively, the $\mathrm{C^{*}}$-algebras $C_{0}(X)$ and $C_{0}(Y)$ are $\sigma$-unital). The following notation is set up to match the hypotheses of [vignati2018rigidity, Theorem 4.3]. We let
@@
The letters $e$ and $o$ are for even and odd partition. Recall ([Black:Operator, II.3.4]) that if $A$ is a $\mathrm{C^{*}}$-algebra, a $\mathrm{C^{*}}$-subalgebra $B\subseteq A$ is hereditary if for all positive $a\leq b$, $b\in B$ implies that $a\in B$, where $a\in A$ is positive if $0\leq a$ (in the abelian setting, $a$ takes only values in $0,\infty)$).
We have the following properties:
@@
Recall, a -homomorphism between nonunital $\mathrm{C^{*}}$-algebras $\psi\colon A\to B$ is nondegenerate ([Black:Operator, II.7.3.8]) if the ideal generated by the image of $\psi$ is dense (in norm topology). This is equivalent to ask that approximate unit for $A$ is sent to an approximate unit of $B$. In the abelian setting, Gel’fand’s duality asserts that nondegenerate -homomorphisms are dual to proper continuous maps. (Recall, a map between topological spaces is proper if inverse images of compact subsets are compact.)
The following is derived from [vignati2018rigidity]. For the reader’s convenience, we shortly sketch its proof, referring heavily to the notation of [vignati2018rigidity].
@@
    For $i\in\{e,o\}$ and $n\neq m$, $\overline{V}_{n}^{i}\cap\overline{V}_{m}^{i}=\emptyset$, and, if $n>\bar{n}$,
@@
    Every compact subset of $X$ intersects only finitely many of the sets $\{V_{n}^{i}\mid n\geq\bar{n},i\in\{e,o\}\}$.
  3. (3)
@@
    $\Psi^{i}=\sum_{n\geq\bar{n}}\psi_{n}^{i}\colon\prod_{n\geq\bar{n}}C_{0}(U_{n}^{i})\to\prod_{n\geq\bar{n}}C_{0}(V_{n}^{i})$
@@
Proof.
Write $X$ as a increasing union of nonempty open sets $X=\bigcup X_{n}$ where for each $n$ we have that $\overline{X_{n}}\subseteq X_{n+1}$, $X_{n+1}\setminus\overline{X_{n}}$ is nonempty, and each $\overline{X_{n}}$ is compact. We furthermore ask that if $K\subseteq X$ is compact then $K\subseteq X_{n}$ for some $n$. This can be done since $X$ is $\sigma$-compact. Writing $B$ for $C_{0}(X)$, we let $e_{n}^{B}\in C_{0}(X)$ be a positive contraction [^3] A contraction in a $\mathrm{C^{*}}$-algebra is an element whose norm is $\leq 1$. such that $e_{n}^{B}[\overline{X_{n}}]=1$ and $e_{n}^{B}[X\setminus X_{n+1}]=0$, so that $e_{n+1}^{B}e_{n}^{B}=e_{n}^{B}$ for all $n$.
We are ready to construct the sets $V_{n}^{i}$. For convenience, we only give the details on how to construct the sets $V_{n}^{e}$; the ‘odd’ sets can be constructed in the same way after appropriately re-indexing.
Let $e_{n}^{A}$ be an approximate identity of positive contractions for $A=C_{0}(Y)$ such that $e_{n}^{A}[\overline{Y}_{n}]=1$ and $e_{n}^{A}(y)\neq 0$ if and only if $y\in Y_{n+1}$, and let $J_{n}=[10n-1,10n+7]$, so that the algebra $A_{J_{n}}$ given in [vignati2018rigidity, Notation 3.3] is precisely $C_{0}(U_{n}^{e})$. Applying the results of §3 and §4 of [vignati2018rigidity], and specifically Lemma 4.2 therein, we get mutually orthogonal [^4] Meaning that if $n\neq m$ then for all $f\in C_{0}(U_{n}^{e})$ and $f^{\prime}\in C_{0}(U_{m}^{e})$ we have $\alpha_{n}(f)\alpha_{m}(f^{\prime})=0$. functions $\alpha_{n}\colon C_{0}(U_{n}^{e})\to C_{0}(X)$ such that the function
$\Gamma^{e}=\sum\alpha_{n}\colon\prod_{n}C_{0}(U_{n}^{e})\to C_{b}(X)$
has the property that for every $f=(f_{n})\in\prod_{n}C_{0}(U_{n}^{e})$ we have a nonmeager ideal $\mathcal{I}_{f}$ such that
$\pi_{X}(\Gamma^{i}(f_{S}))=\Phi(\pi_{Y}(f_{S}))$
for all $S\in\mathcal{I}_{f}$. The mutually orthogonal $\alpha_{n}$s come with two sequences of naturals $j_{n}<k_{n}$ such that the range of $\alpha_{n}$ is contained in $(e_{k_{n}}^{B}-e_{j_{n}}^{B})C_{0}(X)(e_{k_{n}}^{B}-e_{j_{n}}^{B})$ and $\lim_{n}j_{n}=\infty$. This is the key property in showing condition (2), as it gives that if $K\subseteq X$ is compact then the set
$\{n\mid\exists f\in C_{0}(U_{n}^{e})\exists x\in K\,(\alpha_{n}(f)(x)\neq 0)\}$
is finite.
We will now slightly modify the maps $\alpha_{n}$, which themselves are not -homomorphisms. By Ulam stability for approximate maps between abelian $\mathrm{C^{*}}$-algebras (a result proved by Sěrml in [Semrl.USAbel], see also [vignati2018rigidity, §5] or, for details, [farah2022corona, §5.4 and Theorem 5.18]) shows that we can find $\bar{n}$ and mutually orthogonal -homomorphisms $\psi_{n}^{e}\colon C_{0}(U_{n}^{e})\to(e_{k_{n}}^{B}-e_{j_{n}}^{B})C_{0}(X)(e_{k_{n}}^{B}-e_{j_{n}}^{B})$, for $n\geq\bar{n}$ such that $\Psi^{i}=\sum\psi_{n}^{e}$ still lifts elements of $\prod_{n}C_{0}(U_{n}^{e})$ on nonmeager ideals. The sets $V_{n}^{e}$ are now constructed from the maps $\psi_{n}^{e}$ as in [vignati2018rigidity, §4.1] [^5] Injectivity of the maps $\psi_{n}^{e}$ is stated as an assumption there, but it is not used there., by setting
$V_{n}^{e}=\bigcup_{f\in C_{o}(U_{n}^{e})}\mathrm{supp}(\psi^{e}_{n}(f)).$
Since $\lim_{n}j_{n}=\infty$ and $V_{n}^{e}\subseteq X_{k_{n}}\setminus X_{j_{n}}$, if a $K$ is compact then $K$ can intersect at most finitely many of the $V_{n}^{e}$s. This shows condition  (2). The other properties are verified exactly as in §4.1 in [vignati2018rigidity]. This concludes the construction of the sets $V_{n}^{e}$, and the proof of the proposition. ∎
@@
Claim 2.4.
$V_{1}$ is, modulo compact, clopen.
Proof.
We show that $\overline{V}_{1}\setminus V_{1}$ is contained in the compact set $\overline{V}_{\bar{n}}^{e}$. Combining conditions (1) and (2) of Proposition 2.3 gives that
$\overline{V_{1}}=\overline{\bigcup_{n\geq\bar{n}}V_{n}^{e}\cup V_{n}^{o}}=\bigcup_{n\geq\bar{n}}\overline{V_{n}^{e}}\cup\bigcup_{n\geq\bar{n}}\overline{V_{n}^{o}}\subseteq\overline{V_{\bar{n}}^{e}}\cup\bigcup_{n\geq\bar{n}}V_{n}^{e}\cup V_{n}^{o},$
where the equality $\overline{\bigcup_{n\geq\bar{n}}V_{n}^{e}\cup V_{n}^{o}}=\bigcup_{n\geq\bar{n}}\overline{V_{n}^{e}}\cup\bigcup_{n\geq\bar{n}}\overline{V_{n}^{o}}$ is given by the fact that compact subsets of $X$ can only intersect finitely many of the sets $V_{n}^{e}\cup V_{n}^{o}$. This shows that $\overline{V}_{1}\setminus V_{1}$ is contained in the compact set $\overline{V_{\bar{n}}^{e}}$. ∎
(As a side note, a deep understanding of [vignati2018rigidity, §4.1] shows that if the spaces of interest can be written as a union of pairwise disjoint compact open sets, $V_{1}$ can be chosen to be clopen. We shall not need this fact.)
As the continuous proper maps $\gamma_{n}^{i}$ agree on their common domains, the function
@@
is a well-defined continuous function. Furthermore, $\Gamma$ is proper: if $K\subseteq W$ is compact, then there exists $\bar{n}^{\prime}$ such that $K\subseteq\bigcup_{\bar{n}\leq n\leq\bar{n}^{\prime}}U_{n}^{e}\cup U_{n}^{o}$, hence $\Gamma^{-1}[K]$ is a closed subset of $\bigcup_{\bar{n}\leq n\leq\bar{n}^{\prime}}V_{n}^{e}\cup V_{n}^{o}$. As the latter has compact closure, $\Gamma^{-1}[K]$ is compact. Since $\Gamma$ is proper, it induces a -homomorphism
@@
Since every function $f\in C_{b}(Y)$ can be written (modulo $C_{0}(Y)$) as a sum $f=f_{e}+f_{o}$ where, for $i\in\{e,o\}$, $f_{i}\in\prod_{n\geq\bar{n}}C_{0}(U_{n}^{i})$, it is enough to show $\Psi$ lifts $\Phi_{1}$ on $\prod_{n\geq\bar{n}}C_{0}(U_{n}^{i})$, again for $i\in\{e,o\}$. For convenience, let $i=e$, and fix mutually orthogonal positive contractions $g_{n}\in C_{b}(X)$ such that for all $n\geq\bar{n}$
@@
Let $g_{n}=0$ if $n<\bar{n}$. Such functions exist as distinct $V_{n}^{e}$s have disjoint closures. For $S\subseteq{\mathbb{N}}$, let $g_{S}=\sum_{n\in S}g_{n}$. Note that, even if $V_{1}$ is not clopen, it is clopen modulo compact, and therefore the function $\chi_{V_{1}}$ is equal, modulo $C_{0}(X)$, to a continuous function.
@@
(2.1) $\text{if }(x\in V_{1}\setminus K\text{ and }|\tilde{\Phi}(f_{S})(x)|>\varepsilon)\text{ then }x\in\bigcup_{n\in S}V_{n}^{e}.$
This gives that $g_{S}\tilde{\Phi}(f_{S})=_{C_{0}(X)}\chi_{V_{1}}\tilde{\Phi}(f_{S})$. As $S$ is arbitrary, the same reasoning applied to ${\mathbb{N}}\setminus S$, gives that $g_{S}\tilde{\Phi}(f_{{\mathbb{N}}\setminus S})\in C_{0}(X)$. Altogether, we get that